OpenStudy (anonymous):

Which graph shows a function that crosses the x-axis at negative three comma zero and has a rise over run of four-thirds?

OpenStudy (anonymous):

y=mx+b m is your rise over run and b is where you cross your x axis so... m = 4/3 b = -3

OpenStudy (anonymous):

\[y=\frac{ 4 }{ 3 }x-3\]

OpenStudy (anonymous):



OpenStudy (tkhunny):

The problem statement gives an x-intercept, not a y-intercept. y = mx + b We are given the slope, 4/3. y = (4/3)x + b We are given a point, (-3,0) 0 = (4/3)(-3) + b Solve for b.

OpenStudy (anonymous):

You are absolutely right!

OpenStudy (anonymous):

so follw tkhunny's instructions, plot the points and graph the line.